Let y be defined implicitly by the equation ln(3y)=5xy Use implicit differentiation to find the first derivative of y with respect to x


a)dy/dx=


Use implicit differentiation to find the second derivative of y with respect to x


b)d2y/dx2= 


c)Find the point on the curve where d2y/dx2=0




1
Expert's answer
2021-05-07T09:58:36-0400


ln(3y) = 5xy

(a) Differentiating with respect to x,

(dy/dx)*1/y = 5(y+ xdy/dx)

By transposition we have,


dy/dx = (5y2)/(1 - 5xy) ....................(1)



(b) Again differentiating equation (1) with respect to x,



d2y/dx2 = { 10y dy/dx (1- 5xy) + 5(y+ xdy/dx)*(5y2) } / ( 1 - 5xy)2


On further solving we have,


d2y/dx= { 25y3 * (3 -10xy) } / ( 1- 5xy)3 ..........(2)


(c) For the required point we set equation (2) equal to 0,

Hence, d2y/dx= 0

y=0 or xy = 0.3

but y = 0 is not possible

So xy=0.3

substituting the value of xy in the curve we have

y = 1.493

x = 0.201